# Towards Interpretable Deep Generative Models via Causal Representation Learning

### v1

@@ 8 Conclusion and open questions
We have reviewed CRL and its connections to classical statistical methods such as factor analysis and structural equation modeling. CRL is an emerging discipline with roots in machine learning, statistics, and causality, and draws inspiration from a variety of domains to address important practical challenges in the use of generative models in scientific applications where interpretability and causality are key. From a statistical perspective, CRL offers a new twist on classical factor analysis through its principled use of causal semantics (graphs, interventions, etc.).
@@
Although the ideas behind CRL have evolved slowly over many decades, it is still a developing discipline with numerous open questions. We conclude this review with a brief summary of some directions for future inquiry.
Beyond interventions
@@
In this review, we have focused on using interventions for identification. While this is indeed practical in a variety of settings, there is significant interest in generalizing beyond interventions. An emerging paradigm is to use multiple environments (i.e. non-iid data arising from data augmentation, distribution shifts, different locations, times, contexts, etc.) to identify latent structure and/or improve generalization (Peters et al.,, 2016; Huang et al.,, 2020; Perry et al.,, 2022; Mooij et al.,, 2020). Indeed, experimental interventions are a special class of environments. It is an intriguing problem to understand how general environments can enable identifiability when strict interventions are not available; see e.g. Varici et al., (2024); Chen et al., (2024).
Implementation and evaluation
@@
Early work on CRL has mostly focused on the identification problem, however, just as important is the problem of estimation and practical algorithms. Despite substantial advances in practical aspects of training generative models, these advances have not directly translated over to the setting of CRL. Recent work (Section 7) has begun to address these challenges, but no general framework has emerged. Another important problem is evaluating CRL models: Since we can only identify parameters up to indeterminacy transformations, how can we assess the quality of estimators in a principled way? One approach is to use downstream tasks such as out-of-distribution generalization (Lu et al.,, 2021) and extrapolation (Saengkyongam et al.,, 2024).
Statistical theory
@@
There are significant gaps in our statistical understanding of CRL: Finite-sample rates, misspecification, uncertainty quantification, and model selection all remain underexplored. Some recent exceptions in this direction include Acartürk et al., (2024); Buchholz and Schölkopf, (2024), although many problems remain: Since the models are nonparametric, the usual curse of dimensionality applies, but can causal assumptions help mitigate the curse of dimensionality? Alternatively, if we aim only to estimate the graph $\mathsf{G}$ and its parameters $A$, are conventional rates achievable? How can traditional estimation strategies such as variational inference lead to consistent model selection, and what are its tradeoffs? What about robustness and estimation under misspecification? These are just some of the intriguing statistical questions that demand resolution.
Large language models
@@
The development of large language models (LLMs) and more generally, foundation models, has occurred in parallel with CRL. This suggests the tantalizing possibility of developing pre-trained foundation models with causal interpretability. This comes with numerous practical challenges; e.g. connecting the objectives of next-token prediction with causal objectives; principled use of diverse, multi-modal, internet-scale datasets; extracting interpretable concepts; etc. On the latter, recent work has made progress (Rajendran et al.,, 2024). An alternative line of work extracts interpretable concepts post-hoc from pre-trained LLMs using sparse dictionary learning (Bricken et al.,, 2023; Templeton et al.,, 2024). These sparse dictionary learning approaches assume that each token is the linear combination of a small set of concept vectors, selected from a overcomplete concept dictionary. That is, the concept space has a much higher dimension than the data, in contrast to most CRL methods. Investigating these higher-dimensional latent space settings in the context of CRL is an important area for future research.

### v2

@@ 8 Conclusion and open questions
CRL is an emerging discipline with roots in machine learning, statistics, and causality, and draws inspiration from a variety of domains to address important practical challenges in the use of generative models in scientific applications where interpretability and causality are key. From a statistical perspective, CRL offers a new twist on classical factor analysis through its principled use of causal semantics (graphs, interventions, etc.).
@@
Although the ideas behind CRL have evolved slowly over many decades, it is still a developing discipline with numerous open questions. We conclude with a brief summary of some directions for future inquiry.
@@
Beyond interventions. So far we have focused on using interventions for identification. While this is indeed practical in a variety of settings, there is significant interest in generalizing beyond interventions. An emerging paradigm is to use multiple environments (i.e. non-iid data arising from data augmentation, distribution shifts, different locations, times, contexts, etc.) to identify latent structure and/or improve generalization (Peters et al., 2016; Mooij et al., 2020; Huang et al., 2020; Perry et al., 2022). Indeed, experimental interventions are a special type of environment. It is an intriguing problem to understand how general environments can enable identifiability when strict interventions are not available; see e.g. Varici et al. (2024); Chen et al. (2024).
@@
Implementation and evaluation.  Early work on CRL has mostly focused on the identification problem; however, just as important is the problem of estimation and practical algorithms. Despite substantial advances in practical aspects of training generative models, these advances have not directly translated over to the setting of CRL. Recent work (Section 7) has begun to address these challenges, but no general framework has emerged. Another important problem is evaluating CRL models: Since we can only identify parameters up to indeterminacy transformations, how can we assess the quality of estimators in a principled way? One approach is to use downstream tasks such as out-of-distribution generalization (Lu et al., 2021) and extrapolation (Saengkyongam et al., 2024).
Optimization and nonconvexity.  A significant challenge in implementing and evaluating CRL models is that the optimization landscape is highly nonconvex and sensitive to modeling assumptions. This is true even in the linear case, as is well-known for factor models, and these problems are only exacerbated with neural networks. Moreover, in practice, classical nonparametric estimators are replaced with variational approximations that are optimized through stochastic gradient descent. The behaviour of these approximations is an outstanding open problem of significant interest, even for simple one-dimensional generative models (see e.g. Kwon and Chae, 2024). On a positive note, identifiable models often have better-behaved loss surfaces (e.g. well-defined minima); investigating the optimization landscape in identifiable CRL models is an interesting question for further research.
@@
Statistical theory.  There are significant gaps in our statistical understanding of CRL: Finite-sample rates, uncertainty quantification, and misspecification all remain underexplored. Some recent exceptions in this direction include Acartürk et al. (2024); Buchholz and Schölkopf (2024), although many problems remain: Since the models are nonparametric, the usual curse of dimensionality applies, but can causal assumptions help mitigate the curse of dimensionality? Alternatively, if we aim only to estimate the graph $\mathsf{G}$ and its parameters $\bm{A}$, are conventional rates achievable? What about robustness and estimation under misspecification? These are just some of the intriguing statistical questions that demand resolution.
Generalization and transfer.  Transferability and out-of-distribution (OOD) generalization are key motivations for learning causal representations in practice, which are related to recent work on causal invariance (Bühlmann, 2020) and invariant causal prediction (Peters et al., 2016). It remains an important statistical problem to formalize these connections and to understand how and when causal representations lead to genuine improvements in OOD transfer and generalization. Numerous studies point to challenges when using non-causal models (Schott et al., 2021; Montero et al., 2021), and recent work suggests causal models as a potential solution (Kong et al., 2022; Richens and Everitt, 2024).
@@
Large language models.  The development of large language models (LLMs) and more generally, foundation models, has occurred in parallel with CRL. This suggests the tantalizing possibility of developing pre-trained foundation models with causal interpretability. This comes with numerous practical challenges; e.g. connecting the objectives of next-token prediction with causal objectives; principled use of diverse, multi-modal, internet-scale datasets; extracting interpretable concepts; etc. On the latter, an emerging literature explores how to identify interpretable concepts in foundation models (Leemann et al., 2023; Rajendran et al., 2024). This connects to a broader literature on how concepts are embedded in the latent spaces of foundation models via so-called linear representations (e.g. Mikolov et al., 2013; Szegedy et al., 2013). An alternative line of work extracts interpretable concepts post-hoc from pre-trained LLMs using sparse dictionary learning (Bricken et al., 2023; Templeton et al., 2024). These sparse dictionary learning approaches assume that each token is the linear combination of a small set of concept vectors, selected from a overcomplete concept dictionary. That is, the concept space has a much higher dimension than the data, in contrast to most CRL methods. Investigating these higher-dimensional latent space settings in the context of CRL is an important area for future research.
9 Funding
B.A. was supported by NSF IIS-2453378, NSF IIS-1956330, and NIH R01GM140467.